OpenStudy (anonymous):

find the sum 3+10+17+24+...+7003+7010

OpenStudy (anonymous):

i did this ,but im sure im missing something 7010-3/2

OpenStudy (amistre64):

yeah, somethings off with that

OpenStudy (amistre64):

what type of pattern do we have?

OpenStudy (anonymous):

arithmetic sequence

OpenStudy (amistre64):

good, then we should be able to add the first and last as part of the process. knowing how many terns there are tho is also needed

OpenStudy (amistre64):

\[\frac{n}{2}(first+last)\]

OpenStudy (anonymous):

so then it would be (7010)/2(3+7010) like this

OpenStudy (amistre64):

there are not 7010 terms, but you are on the right track

OpenStudy (amistre64):

what is our common difference? what is added to get a new term?

OpenStudy (anonymous):

there is a finite number which stops at 7010 so should i apply the formula an=a1+(n-1)d

OpenStudy (amistre64):

you should yes

OpenStudy (anonymous):

but i don't know the common difference it goes 3 to 7 to 7 again it's not constant

OpenStudy (amistre64):

3+10+17+24+...+7003+7010 7 7 7 7

OpenStudy (anonymous):

ohhh oops lol

OpenStudy (anonymous):

i didn't follow clearly no wonder

OpenStudy (amistre64):

on a small scale, lets say our difference is 4: 2,6,10,14 we have 4 terms we can adjust it by subtracting 2 from it all 0,4,8,12, now divide by 4 0,1,2,3 ... the number of terms is 1 more than the last adjusted term

OpenStudy (amistre64):

n = (7010-3)/7 + 1

OpenStudy (amistre64):

i spose we could have done the same with the rule: an = ao + (n-1)d an - ao = (n-1)d (an - ao)/d = n-1 (an - ao)/d + 1 = n

OpenStudy (anonymous):

an=3+(7010-1)7

OpenStudy (amistre64):

n-1, not 7010-1 you are trying to find the value of n

OpenStudy (anonymous):

the an term? would be 7010 is that right

OpenStudy (anonymous):

ohh ok

OpenStudy (amistre64):

7010 = 3 + 7(n-1) yes, solve for n

OpenStudy (anonymous):

makes sense now thanks for the clarification

OpenStudy (amistre64):

good luck ;)